# Lagrangian extensions and left symmetric structures on the four-dimensional real Lie superalgebras

### v1

@@ 4. The 4-dimensional Lie superalgebras as Lagrangian extensions
    The Lie superalgebra $(D^{10}_{0})^{1}$: In the case where the form is odd, this Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{4},e_{4}]_{\mathfrak{h}}=e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=-\Pi(e_{4}^{*})|e_{3}=\Pi(e_{1}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
@@
    In the case where the form is even, this Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{4},e_{4}]_{\mathfrak{h}}=e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=2e_{1}^{*}|e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
@@
    The Lie superalgebra $(D^{10}_{0})^{2}$: In the case where the form is odd, this Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{4},e_{4}]_{\mathfrak{h}}=-e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=-\Pi(e_{4}^{*})|e_{3}=\Pi(e_{1}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
@@
    In the case where the form is even, this Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{4},e_{4}]_{\mathfrak{h}}=-e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=2e_{1}^{*}|e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
@@
    The connection $\nabla$ is given by:
@@
    The Lie superalgebra $(C^{1}_{1/2}+A)$: This Lie superalgebra does not admit a 2-dimensional Lagrangian ideal. Indeed, let us suppose there is one, denoted by $\mathfrak{a}=\mathrm{Span}\{X,Y\}$. Let us write $X=x_{1}e_{1}+x_{2}e_{2}+x_{3}e_{3}+x_{4}e_{4}$ and $Y=y_{1}e_{1}+y_{2}e_{2}+y_{3}e_{3}+y_{4}e_{4}$. We compute the brackets $[e_{i},X]$ for all $i\in\{1,2,3,4\}$. We get:
@@
    The Lie superalgebra $(C^{3}+A)$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{2}|e_{4}\}$ with the bracket $[e_{4},e_{4}]_{\mathfrak{h}}=e_{2}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{1}=2e_{2}^{*}\,|\,e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
@@
    The Lie superalgebra $D^{1}$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
@@
    The Lie superalgebra $D^{2}_{q},\text{ with }q\not=-1,0$: This Lie superalgebra is not quasi-Frobenuis.
@@
    The Lie superalgebra $D^{2}_{-1}$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
@@
    The Lie superalgebra $D^{11}_{pq}\text{ with }(p,q)=(-1,-1)$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
@@
    The Lie superalgebra $D^{11}_{pq}\text{ with }0<|p|\leq 1,p=-q$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
@@
    The Lie superalgebra $D^{11}_{pq}\text{ with }0<|p|<1,q=-1$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
@@
    The Lie superalgebra $D^{12}$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
@@
    The Lie superalgebra $D^{13}_{-1}$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
@@
    The Lie superalgebra $D^{15}$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
@@
    The Lie superalgebra $D^{16}$: This Lie superalgebra is in fact not quasi-Frobenuis.
@@
    The Lie superalgebra $(D^{3}_{-1/2,-1/2})$: This Lie superalgebra doesn’t admit a non-degenerate homogeneous closed form.
@@ 5. Left-symmetric structures on the 4-dimensional real Lie superalgebras
In the list of 4-dimensional real Lie superalgebras, we will list left-symmetric structures that are compatible with the Lie structure. With the exception of $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$, all of these structures are Novikov. Moreover, we show that each of these Lie superalgebras admit a Balinsky-Novikov structure. In line with standard practice, only non-zero products are displayed.
@@
    The Lie superalgebra $D^{5}$: A Novikov structure is given by:
@@
    Additionally, this structure is also Balinsky-Novikov. As a matter of fact, the non-zero associators are given by:
@@
    The Lie superalgebra $D^{6}$: A Novikov structure is given by:
@@
    The Lie superalgebra $D^{7}_{pq}$ with ($p\geq q,\;pq\not=0$): A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
@@
    The Lie superalgebra $D^{8}_{p}$ with $p\not=0$: A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
@@
    The Lie superalgebra $D^{9}_{pq}$ with $q>0$: A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
@@
    A Balinsky-Novikov structure is given by (where $\gamma\in\mathbb{R}$):
@@
    The Lie superalgebra $D^{10}_{q}:$ A Novikov structure is given by (where $\gamma\in\mathbb{R}$):
@@ A Balinsky-Novikov structure is given by
[…]
@@
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
@@
[…]
@@
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
@@
[…]
@@
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
@@
    The Lie superalgebra $(D^{7}_{1-p,p})$ with $p\leq\frac{1}{2}:$ A Novikov structure is given by (where $\lambda\in\mathbb{R}$):
@@
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
@@
    The Lie superalgebra $(D^{8}_{1/2}):$ A Novikov structure is given by (where $\lambda\in\mathbb{R}$):
@@
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=\gamma e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=(1+\gamma)e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=e_{3}+\frac{1}{2}(1+\gamma)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=\gamma e_{2},&e_{3}\cdot e_{1}=\gamma e_{3},&e_{4}\cdot e_{1}=\gamma e_{4},&e_{4}\cdot e_{4}=e_{2}.\\[5.69054pt] \end{array}$
@@
    The Lie superalgebra $(D^{9}_{1/2,p})$ with $p>0:$ A Novikov structure is given by (where $\lambda\in\mathbb{R}$):
@@
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
@@
    The Lie superalgebra $(D^{10}_{0})^{1}:$ An LSSA structure is given by (where $\gamma\in\mathbb{R}$):
@@
    However, there is a Balinsky-Novikov structure given by (where $\lambda,\gamma\in\mathbb{R}$):
@@
    The Lie superalgebra $(D^{10}_{0})^{2}:$ AAn LSSA structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
@@
    However, there is a Balinsky-Novikov structure given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=(\gamma-1)e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},&e_{1}\cdot e_{4}=-\lambda e_{3}+\frac{1}{2}(\gamma-1)e_{4},\\[5.69054pt] e_{2}\cdot e_{1}=(\gamma-1)e_{2},&e_{2}\cdot e_{4}=(1-\gamma)e_{3},&e_{3}\cdot e_{1}=(\gamma-1)e_{3},&e_{3}\cdot e_{4}=\frac{1}{2}e_{2},\\[5.69054pt] e_{4}\cdot e_{1}=-2\lambda e_{3}+(\gamma-1)e_{4},&e_{4}\cdot e_{2}=-2\gamma e_{3}&e_{4}\cdot e_{3}=\frac{1}{2}e_{2},&e_{4}\cdot e_{4}=-e_{1}.\end{array}$
@@
    The Lie superalgebra $(2A_{1,1}+2A)^{1}:$ A Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R})$:
@@
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
@@
    The Lie superalgebra $(2A_{1,1}+2A)^{2}:$ A Novikov structure is given by (where $\gamma\in{\mathbb{R}})$::
    $\begin{array}[]{l}e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\quad e_{3}\cdot e_{4}=\gamma e_{1},\quad e_{4}\cdot e_{3}=(1-\gamma_{)}e_{1},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure is given by (where $\lambda,\gamma\in\mathbb{R}$):
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=2\gamma e_{1},&e_{1}\cdot e_{2}=2\gamma e_{1},&e_{1}\cdot e_{3}=\gamma e_{3},&e_{1}\cdot e_{4}=\gamma e_{3},\\[5.69054pt] e_{2}\cdot e_{1}=2\gamma e_{1},&e_{2}\cdot e_{2}=-2(\lambda-\gamma)e_{1}+2\lambda e_{2},&e_{2}\cdot e_{3}=\gamma e_{3},&e_{2}\cdot e_{4}=(\gamma-\lambda)e_{3}+\lambda e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=2\gamma e_{3}&e_{3}\cdot e_{2}=2\gamma e_{3},&e_{3}\cdot e_{3}=e_{1}&e_{3}\cdot e_{4}=e_{1},\\[5.69054pt] e_{4}\cdot e_{1}=2\gamma e_{3}&e_{4}\cdot e_{2}=-2(\lambda-\gamma)e_{3}+2\lambda e_{4}&e_{4}\cdot e_{3}=e_{1}&e_{4}\cdot e_{4}=e_{2}.\end{array}$

### v2

@@ 4. The 4-dimensional Lie superalgebras as Lagrangian extensions
    The Lie superalgebra $(D^{10}_{0})^{1}$: In the case where the form is odd, this Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{4},e_{4}]_{\mathfrak{h}}=e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=-\Pi(e_{4}^{*})|e_{3}=\Pi(e_{1}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by
@@
    In the case where the form is even, this Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{4},e_{4}]_{\mathfrak{h}}=e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=2e_{1}^{*}|e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by
@@
    The Lie superalgebra $(D^{10}_{0})^{2}$: In the case where the form is odd, this Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{4},e_{4}]_{\mathfrak{h}}=-e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=-\Pi(e_{4}^{*})|e_{3}=\Pi(e_{1}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by
@@
    In the case where the form is even, this Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{4},e_{4}]_{\mathfrak{h}}=-e_{1}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=2e_{1}^{*}|e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by
@@
    The connection $\nabla$ is given by
@@
    The Lie superalgebra $(C^{1}_{1/2}+A)$: This Lie superalgebra does not admit a Lagrangian ideal of dimension 2. Indeed, let us suppose there is one, denoted by $\mathfrak{a}=\mathrm{Span}\{X,Y\}$. Let us write $X=x_{1}e_{1}+x_{2}e_{2}+x_{3}e_{3}+x_{4}e_{4}$ and $Y=y_{1}e_{1}+y_{2}e_{2}+y_{3}e_{3}+y_{4}e_{4}$. We compute the brackets $e_{i},X]$, for all $i\in\{1,2,3,4\}$. We obtain
@@
    The Lie superalgebra $(C^{3}+A)$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{2}|e_{4}\}$ with the bracket $e_{4},e_{4}]_{\mathfrak{h}}=e_{2}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{1}=2e_{2}^{*}\,|\,e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by
@@
    The Lie superalgebra $D^{1}$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
@@
    The Lie superalgebra $D^{2}_{q},\text{ with }q\not=-1,0$: This Lie superalgebra is indeed not quasi-Frobenuis.
@@
    The Lie superalgebra $D^{2}_{-1}$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
@@
    The Lie superalgebra $D^{11}_{pq}\text{ with }(p,q)=(-1,-1)$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
@@
    The Lie superalgebra $D^{11}_{pq}\text{ with }0<|p|\leq 1,p=-q$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
@@
    The Lie superalgebra $D^{11}_{pq}\text{ with }0<|p|<1,q=-1$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
@@
    The Lie superalgebra $D^{12}$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
@@
    The Lie superalgebra $D^{13}_{-1}$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
@@
    The Lie superalgebra $D^{15}$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
@@
    The Lie superalgebra $D^{16}$: This Lie superalgebra is not quasi-Frobenuis.
@@
    The Lie superalgebra $(D^{3}_{-1/2,-1/2})$: This Lie superalgebra does not admit a non-degenerate homogeneous closed form.
@@
Tables 15 below summarize the outcome of the classification. Among the Lie superalgebras considered, 22 do not admit a quasi-Frobenius structure, whereas 9 admit a nonhomogeneous structure. Of the quasi-Frobenius cases, 4 do not arise as Lagrangian extensions. Moreover, 8 are $T^{*}$-extensions and 10 are $\Pi T^{*}$-extensions. It is worth mentioning that three Lie superalgebras admit both a closed orthosymplectic and periplectic form, namely $(D_{0}^{10})^{1}$, $(D_{0}^{10})^{2}$, and $D_{pq}^{7}$ with (p,q)=(1,1)(p,q)=(1,-1).
Lie superalgebras without quasi-Frobenius structures
$(D_{1/2\;1/2}^{7})^{3}$ $(D^{8}_{1/2})$ $(2A_{1,1}+2A)^{1}$
$(2A_{1,1}+2A)^{2}$ $D^{7}_{pq}$ $(A_{1,1}+3A)^{1}$
$(A_{1,1}+3A)^{2}$ $(D^{2}_{-1/2})^{1}$ $(D^{2}_{-1/2})^{2}$
$(D^{3}_{p,-1/2}),\,p\not=0$ $(A_{3,1}+A)$ $D^{16}$
$D^{14}_{pq},\,p\not=0,q\geq 0$ $D^{13}_{p},\,p\not=0$ $D^{3}_{pq}$
$D^{11}_{pq},\,0<|p|\leq|q|\leq 1$ $D^{2}_{q},\,q\not=-1,0$ $D^{9}_{pq},\,q>0$
$D^{8}_{p},\,p\not=0$ $(C_{0}^{5}+A)$ $(C^{2}_{-1}+A)$
$(2A_{1,1}+2A)^{4}_{p}$
Table 1. No quasi-Frobenius structures
Lie superalgebras with a nonhomogeneous quasi-Frobenius structure
$(D^{3}_{-1/2\;1/2})$ $D^{15}$ $D^{13}_{-1}$
$D^{12}$ $D^{11}_{pq},\,0<|p|<1,\,q=-1$ $D^{11}_{pq},\,0<|p|<1,\,p=-q$
$D^{11}_{pq},\,p=q=-1$ $D^{2}_{-1}$ $D^{1}$.
Table 2. Nonhomogeneous
Quasi-Frobenuis Lie superalgebras that are not Lagrangian extensions
$(C_{1/2}^{1}+A)$ $(D^{9}_{1/2\;p}),\,p>0$ $(D^{7}_{1/2\;1/2})^{1}$
$D^{6}$
Table 3. Not Lagrangian extensions
Quasi-Frobenuis Lie superalgebras that are $T^{*}$-extensions
$(C^{3}+A)$ $(C_{1}^{1}+A)$ $(D_{0}^{10})^{1},\,(|\omega|=\bar{0})$
$(D_{0}^{10})^{2},\,(|\omega|=\bar{0})$ $D^{7}_{1-p,p},\,p\leq\frac{1}{2}$ $(D^{7}_{1/2\;1/2})^{2}$
$D^{7}_{pq},\,p=-q,q\not=0,-1$ $D^{7}_{pq},\,p=-q=1$ $(|\omega|=\bar{0})$
Table 4. $T^{*}$-extensions
Quasi-Frobenuis Lie superalgebras that are $\Pi T^{*}$-extensions
$(2A_{1,1}+2A)^{3}_{p},\,p=\frac{1}{2}$ $(D_{0}^{10})^{2},\,(|\omega|=\bar{1})$ $(D_{0}^{10})^{1},\,(|\omega|=\bar{1})$
$D^{10}_{q},\,q\not=-1$ $D^{6}$ $D^{7}_{-1,q},\,q\leq 1$
$D^{7}_{p,q},\,p=-q=1,\,(|\omega|=\bar{1})$ $D^{5}$ $D^{8}_{-1}$
$D^{7}_{p,q},\,p=-q=1$
Table 5. $\Pi T^{*}$-extensions
@@ 5. Left-symmetric structures on the 4-dimensional real Lie superalgebras
In the list of indecomposable 4-dimensional real Lie superalgebras, we will list left-symmetric structures that are compatible with the Lie structure. With the exception of $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$, all of these structures are Novikov. Moreover, we show that each of these Lie superalgebras admit a Balinsky-Novikov structure. In line with standard practice, only non-zero products are displayed.
@@
    The Lie superalgebra $D^{5}$: A Novikov structure is given by
@@
    Additionally, this structure is also Balinsky-Novikov. As a matter of fact, the non-zero associators are given by
@@
    The Lie superalgebra $D^{6}$: A Novikov structure is given by
@@
    The Lie superalgebra $D^{7}_{pq}$ with ($p\geq q,\;pq\not=0$): A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $D^{8}_{p}$ with $p\not=0$: A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $D^{9}_{pq}$ with $q>0$: A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $D^{10}_{q}:$ A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
@@ A Balinsky-Novikov structure is given by
    The Lie superalgebra $(D^{7}_{1/2\;1/2})^{1}:$ A Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $(D^{7}_{1/2\;1/2})^{2}:$ A Novikov structure, depending on a parameter $\lambda\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $(D^{7}_{1/2\;1/2})^{3}:$ A Novikov structure, depending on a parameter $\lambda\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $(D^{7}_{1-p,p})$ with $p\leq\frac{1}{2}:$ A Novikov structure, depending on a parameter $\lambda\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $(D^{8}_{1/2}):$ A Novikov structure, depending on a parameter $\lambda\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
[…]
@@
    The Lie superalgebra $(D^{9}_{1/2,p})$ with $p>0:$ A Novikov structure, depending on a parameter $\lambda\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $(D^{10}_{0})^{1}:$ An LSSA structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
@@
    However, there is a Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $(D^{10}_{0})^{2}:$ An LSSA structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    However, there is a Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by given by
    $\begin{array}[]{lll}e_{1}\cdot e_{1}=(\gamma-1)e_{1}+\lambda e_{2},&e_{1}\cdot e_{2}=\gamma e_{2},&e_{1}\cdot e_{3}=\frac{1}{2}(1+\gamma)e_{3},\\[5.69054pt] e_{1}\cdot e_{4}=-\lambda e_{3}+\frac{1}{2}(\gamma-1)e_{4},&e_{2}\cdot e_{1}=(\gamma-1)e_{2},&e_{2}\cdot e_{4}=(1-\gamma)e_{3},\\[5.69054pt] e_{3}\cdot e_{1}=(\gamma-1)e_{3},&e_{3}\cdot e_{4}=\frac{1}{2}e_{2},&e_{4}\cdot e_{1}=-2\lambda e_{3}+(\gamma-1)e_{4},\\[5.69054pt] e_{4}\cdot e_{2}=-2\gamma e_{3}&e_{4}\cdot e_{3}=\frac{1}{2}e_{2},&e_{4}\cdot e_{4}=-e_{1}.\end{array}$
@@
    The Lie superalgebra $(2A_{1,1}+2A)^{1}:$ A Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
@@
    The Lie superalgebra $(2A_{1,1}+2A)^{2}:$ A Novikov structure, depending on a parameter $\gamma\in\mathbb{R}$, is given by
    $\begin{array}[]{l}e_{3}\cdot e_{3}=\frac{1}{2}e_{1},\quad e_{3}\cdot e_{4}=\gamma e_{1},\quad e_{4}\cdot e_{3}=(1-\gamma)e_{1},\quad e_{4}\cdot e_{4}=\frac{1}{2}e_{2}.\end{array}$
    A Balinsky-Novikov structure, depending on parameters $\lambda,\gamma\in\mathbb{R}$, is given by
    $\begin{array}[]{llll}e_{1}\cdot e_{1}=2\gamma e_{1},&e_{1}\cdot e_{2}=2\gamma e_{1},&e_{1}\cdot e_{3}=\gamma e_{3},&e_{1}\cdot e_{4}=\gamma e_{3},\\[5.69054pt] e_{2}\cdot e_{1}=2\gamma e_{1},&e_{2}\cdot e_{2}=-2(\lambda-\gamma)e_{1}+2\lambda e_{2},&e_{2}\cdot e_{3}=\gamma e_{3},&e_{2}\cdot e_{4}=(\gamma-\lambda)e_{3}+\lambda e_{4},\\[5.69054pt] e_{3}\cdot e_{1}=2\gamma e_{3},&e_{3}\cdot e_{2}=2\gamma e_{3},&e_{3}\cdot e_{3}=e_{1},&e_{3}\cdot e_{4}=e_{1},\\[5.69054pt] e_{4}\cdot e_{1}=2\gamma e_{3},&e_{4}\cdot e_{2}=-2(\lambda-\gamma)e_{3}+2\lambda e_{4},&e_{4}\cdot e_{3}=e_{1},&e_{4}\cdot e_{4}=e_{2}.\end{array}$